# SymBoltz.jl: a symbolic-numeric, approximation-free and differentiable linear Einstein-Boltzmann solver

### v2

@@ 1.3 Differentiation methods
of (many) elementary operations $\boldsymbol{f}_{i}:\mathbb{R}^{a_{i}}\rightarrow\mathbb{R}^{b_{i}}$ (think of $\boldsymbol{f}_{i}$ as the $i$-th line of code). It then numerically evaluates the chain rule
$\boldsymbol{J}=\boldsymbol{J}_{N}\cdot\boldsymbol{J}_{N-1}\cdots\boldsymbol{J}_{2}\cdot\boldsymbol{J}_{1}.$ (2)
through the Jacobian $(\boldsymbol{J}_{n})_{ij}=\partial f_{n,i}/\partial f_{n-1,j}$ of every elementary operation to accumulate the derivative of the entire program. This is numerically exact, free of precision parameters, usually requires fewer operations than finite differences, but is perhaps less intuitive, harder to implement and needs the source code of the program (1) to interpret it in the non-standard way (2).
Notably, while the function (1) must be evaluated inside-to-outside (right-to-left), the chain rule (2) is an associative matrix product that can be evaluated in any order. This generally changes the number of operations and is a more open-ended computational problem. Forward-mode automatic differentiation seeds $\boldsymbol{J}_{1}=\boldsymbol{1}$ (the derivative of the input with respect to itself) and multiplies $\boldsymbol{J}_{N}(\boldsymbol{J}_{N-1}(\cdots(\boldsymbol{J}_{2}(\boldsymbol{J}_{1}))))$ by ?pushing? every column of $\boldsymbol{J}_{1}$ forward through the product in the same evaluation order as $\boldsymbol{f}$. Reverse-mode first computes $\boldsymbol{f}$ in a forward pass, then seeds $\boldsymbol{J}_{N}=\boldsymbol{1}$ (the derivative of the output with respect to itself) and multiplies $((((\boldsymbol{J}_{N})\boldsymbol{J}_{N-1})\cdots)\boldsymbol{J}_{2})\boldsymbol{J}_{1}$ by ?pulling? every row of $\boldsymbol{J}_{N}$ backwards through the product. This usually makes forward-mode faster when $\boldsymbol{f}:\mathbb{R}^{a}\rightarrow\mathbb{R}^{b}$ has more outputs $b\gg a$, and reverse-mode better when there are more inputs $a\gg b$.
In practice, both modes are implemented using techniques like operator overloading or source code transformation. The former specializes every function on a particular ?dual number? that computes both the value and derivative of the function (e.g. Revels et al., 2016). The latter analyzes the source code for $\boldsymbol{f}$ and transforms it into another code that computes $\boldsymbol{J}$. In any case, automatic differentiation requires access to the source code!
@@
Figure 1: SymBoltz represents cosmological models with symbolic equations grouped in physical components for the metric, gravity and particle species. This is compiled to a numerical problem that splits equations into background and perturbation stages and generates fast code for ODE functions $\boldsymbol{f}$ and Jacobians $\boldsymbol{J}$. The problem is then solved and the result stored in a solution object that gives access to any variable $S$ in the model.
@@ 2 Code architecture and main features
SymBoltz is designed around three main features. In short, it has a symbolic-numeric abstraction interface where users enter high-level symbolic equations, and automatically compiles them to fast low-level numerical functions. It cures stiffness with implicit ODE solvers, and uses no approximation schemes to keep models simple, elegant and extensible. It is differentiable, so one can get accurate derivatives of any output with respect to any input. This provides rapid model prototyping, helpful abstractions and automates tasks one must do manually when modifying other codes. SymBoltz encourages interactive use and pursues a modular design that lets users integrate what they need from the package into their own applications. The end goal is to prioritize the modeler, who should be able to just write down their equations while SymBoltz handles modeling chores.
The code is written in the Julia programming language (Bezanson et al., 2017), with a rich ecosystem of scientific packages and an aim to resolve the two-language problem. The symbolic-numeric interface is built on ModelingToolkit.jl (Ma et al., 2022) and Symbolics.jl (Gowda et al., 2022). The compiled functions are integrated by implicit ODE solvers in OrdinaryDiffEq.jl (Rackauckas and Nie, 2017). Automatic differentiation works through ForwardDiff.jl (Revels et al., 2016).
The next subsections describe SymBoltz’ three main features in more detail. Other implementation details are given in appendix A. This paper is as of SymBoltz version 0.14.1. Please see the package documentation for definitive up-to-date information.
@@ 2.1 Symbolic-numeric interface
The core of SymBoltz is built around a symbolic-numeric interface illustrated in fig. 1. Variables and equations are specified in a high-level user-friendly symbolic modeling language, then compiled down to low-level numerical code that is integrated by ODE solvers. With knowledge of the symbolic equations, SymBoltz can analyze their structure programmatically to ease model specification, automate related boilerplate tasks and generate fast and stable code that avoids approximations.
We think three properties motivate some kind of symbolic abstraction layer around the Einstein-Boltzmann equations.
@@
First, the equations are most easily written as one large ODE, but optimally solved as several ODEs in stages like the background, perturbations and line-of-sight integration. Mathematically, the Jacobian matrix of this single ODE encodes which variables are independent and belong to which stages. It is easiest for modelers to specify equations in one common system, where the perturbations can refer to variables in the background, for example. The separation into computational stages can be automated.
@@
Second, solving each stage is hard due to stiffness, performance requirements and its dependence on previous stages. To overcome this, it is helpful to automatically generate ODE code that is optimal in speed, accuracy and stability, generate the ODE Jacobian needed by implicit solvers and interpolate variables from previous stages. Modelers should not worry about this.
@@
Third, the equations are often modified because the true cosmology is unknown. This can be made easier by organizing equations in a way that reflects the physical structure of the model, so it is easy to replace a related subset of equations with another without duplicating the entire model. This could be slow to do at the numerical level, but can be done with no runtime penalty in an intermediate symbolic representation of the equations.
@@
Inspired by this, SymBoltz inverts the traditional layout of Boltzmann solvers. While most codes are built as pipelines that follow the background, perturbations and other stages of the equations, SymBoltz is primarily structured around the physical components that they consists of. Related variables and equations are grouped in distinct components for the metric, gravitational theory, photons, baryons, dark matter, dark energy, neutrinos and other species (see fig. 1). This isolates everything related to one submodel to one place. Any set of such submodels are joined into a full cosmological model, like $\mathrm{\Lambda CDM}$. Interactions (e.g. Compton scattering or sourcing of gravity) are equations between components. Adding new submodels is easy, and SymBoltz is largely devoted to simply building a well-organized library of them.
@@
SymBoltz then takes a full cosmological model, separates it into stages and generates numerical code to solve each stage. This preprocessing does not slow the code at runtime; to the contrary the symbolic equations give extra information like the Jacobian that enhances speed and stability. Modifications automatically enjoy these benefits without extra user input.
@@
The modular component-based structure scales well in model space. For example, modified gravity or dark energy models can be written down as self-contained submodels and combined with other components to create many different extended cosmological models. The generic symbolic framework can also be used to build reduced models with non-interacting radiation, matter and dark energy species. When exploring modified gravity theories, it can be helpful to test such toy models to understand how gravity responds to pure fluids without coupling to baryons and photons.
@@
In contrast, the monolithic layout of traditional codes scatter changes related to one submodel across different modules for input/output, the background, perturbations and so on. As more and more submodels are added, the code is increasingly intertwined and fragmented. They are also not flexible enough to test simpler toy models, as baryons and photons are hardcoded as fundamental to the structure of the code. This design fits the computational structure better than the physical one. It can result in an overwhelming code that becomes hard to read and modify, and whose complexity scales poorly in model space.
@@
As an alternative to component-based modeling, SymBoltz also includes a version of the default $\mathrm{\Lambda CDM}$ model with all equations packed into one big ?unstructured? system. This makes it trivial to change anything in the equations, but sacrifices modularity. The symbolic interface makes it very compact: SymBoltz defines a full $\mathrm{\Lambda CDM}$ model in just 277 straightforward lines of code [^8] https://hersle.github.io/SymBoltz.jl/stable/unstructured/, while the equivalent code to browse in CLASS is spread over 10 files with 27721 lines [^9] Counting input.{h,c}, background.{h,c}, thermodynamics.{h,c}, perturbations.{h,c} and wrap_recfast.{h,c} with wc -l..
PyCosmo’s symbolic interface also generates code for ODEs and sparse Jacobians, but does not focus on component-based modeling or automating tasks like stage separation, for example.
@@
SymBoltz’ symbolic-numeric interface simplifies solving and modifying the Einstein-Boltzmann equations. It aims to maximize speed, stability and convenience with minimal user input. The next sections go into detail on how this is done. Section 3.2 shows a concrete example and some of its features in action.
@@ 2.1.1 Automatic numerical code generation
SymBoltz automatically compiles symbolic equations to numerical code for ordinary differential equations (ODEs)
$\frac{\mathrm{d}\boldsymbol{u}}{\mathrm{d}\tau}=\boldsymbol{f}(\tau,\boldsymbol{u},\boldsymbol{p}).$ (3)
The generated code is fast and prevents users unfamiliar with Julia or SymBoltz from writing slow code. If necessary, one can escape the standard code generation and call arbitrary numerical functions, for example to solve a nonlinear equation for the minimum of a potential or interpolate tabulated data. The code generation handles tasks like allocating indices for each ODE state $u_{i}$, and performs optimizations such as common subexpression elimination (see section 3.2). This is helpful as Boltzmann solvers tend to have big $\boldsymbol{f}$ and ODE solvers call $\boldsymbol{f}$ many times.
@@ 2.1.2 Automatic handling of observed variables
A general ODE (3) has two types of variables: ?unknowns? $\boldsymbol{u}(\tau)$ are integrated with respect to time, while ?observed? variables are any functions of the unknowns. The Einstein-Boltzmann equations are usually written with many observed variables.
For example, consider the metric and gravity equations in sections A.1 and A.2 sourced by some arbitrary $\rho(\tau)$, $\delta\rho(\tau,k)$ and $\Pi(\tau,k)$. Here $a(\tau)$ and $\Phi(\tau,k)$ are the only unknown (differential) variables that the ODE is solved for, while one can ?observe? $z(\tau)$, $\mathscr{H}(\tau)$, $H(\tau)$ and $\Psi(\tau,k)$ from the unknowns. In addition, observed derivatives like $\mathscr{H}^{\prime}$ or $\Psi^{\prime}$ are not always trivial to express, but SymBoltz automatically expands them using the definitions of $\mathscr{H}$ or $\Psi$. Of course, one can eliminate all observeds by explicitly inserting them into the equations for the unknowns. But this reduces readability, as observed variables are helpful intermediate definitions that break up the equations, and one may want to extract them from the solution as well. Furthermore, modified models can change the sets of unknown and observed variables (e.g. modified gravity can change the constraint equation for $\Psi$ into a differential equation). It is easier to compose models when variables are not hardcoded as either unknown or observed.
SymBoltz reads in a full system of equations like that defined by appendix A and separates variables into unknowns and observeds. It then generates numerical code for solving the ODE for the unknowns (see section 2.1.1), but can automatically recompute any observed variables from the solution of the unknowns.
The bottom line is that users can easily use any variable anywhere by just referring to it, whether it is unknown or observed. In other solvers one may have to look up the expression for observed variables and recompute them manually. This is tedious, error-prone and can tempt users to make simplifying assumptions (e.g. approximating $\Psi\approx\Phi$ without anisotropic stress).
@@ 2.1.3 Automatic stage separation and splining of unknowns
In principle, one can solve the Einstein-Boltzmann equations by integrating the entire system (e.g. background and perturbations) at once. However, the system can be broken down into sequential computational stages that each depend only on those before it. To alleviate stiffness in each stage, separate concerns by solving every equation only once (e.g. avoid recomputing the background for every perturbation mode) and to improve performance, all Boltzmann codes solve the system stage-by-stage and spline variables from one stage as input to the next.
@@
To illustrate, again consider the general relativistic equations in section A.2 sourced by some known $\rho(\tau)$, $\delta\rho(\tau,k)$ and $\Pi(\tau,k)$. Clearly $\Phi(\tau,k)$ and $\Psi(\tau,k)$ depend on $a(\tau)$, but $a(\tau)$ does not depend on $\Phi(\tau,k)$ or $\Psi(\tau,k)$, reflecting the perturbative nature of the problem. One can first solve for only $a(\tau)$ in the ?background?, then spline and look up $a(\tau)$ to solve for $\Phi(\tau,k)$ and $\Psi(\tau,k)$ in the ?perturbations?, instead of solving for all three together and repeatedly solve for $a(\tau)$ for every $k$.
@@
SymBoltz uses the same stage separation strategy as any other Boltzmann code, but automates it with its knowledge of the symbolic equations. First, all equations are split into background and perturbation stages. Cubic Hermite splines are then constructed for all background unknowns (like $a(\tau)$) and passed on to the perturbations. This spline type is optimal for interpolating ODEs, as Hermite splines take both $\boldsymbol{u}(\tau)$ and $\boldsymbol{u}^{\prime}(\tau)$ into account for better accuracy, and $\boldsymbol{u}^{\prime}(\tau)$ is known analytically from $\boldsymbol{u}(\tau)$ and definition (3). In contrast, observed variables (like $\mathscr{H}(\tau)$) are computed from the (splined) unknowns because their derivatives are not known directly, and splining them is less accurate. This makes any background variable available in the perturbations ?for free?. All stages of the Einstein-Boltzmann equations are solved as if it is written as one combined system of equations, while splining and stage separation are done automatically under the hood.
@@
The separation into background and perturbation stages is guaranteed and just reflects the perturbative structure of the linearized Einstein-Boltzmann equations, where each order depends only on those before it. However, they can often be broken further down: most thermodynamics (recombination) models can be separated from the background, and some variables have integral solutions (e.g. the optical depth $\kappa(\tau)=\int_{\tau_{0}}^{\tau}\kappa^{\prime}(\bar{\tau})\mathrm{d}\bar{\tau}$ or line-of-sight integration) that can be computed in isolation after solving the differential equations. In the future, SymBoltz’ background-perturbations separation could be extended to automatically split all smaller stages by symbolically inspecting the dependencies between variables in the equations.

### v3

@@ 1 Introduction
of elementary operations $\boldsymbol{f}_{i}:\mathbb{R}^{m_{i}}\rightarrow\mathbb{R}^{n_{i}}$ (think of $\boldsymbol{f}_{i}$ as the $i$-th line of code). Then it numerically evaluates the chain rule
$\boldsymbol{J}=\boldsymbol{J}_{N}\cdot\boldsymbol{J}_{N-1}\cdot\cdots\cdot\boldsymbol{J}_{2}\cdot\boldsymbol{J}_{1}$ (2)
through the Jacobian $(\boldsymbol{J}_{n})_{ij}=\partial f_{n,i}/\partial f_{n-1,j}$ of every operation to accumulate the derivative of the entire program. This is numerically exact and free of precision parameters, while usually requiring fewer operations than finite differences. But it is perhaps less intuitive, harder to implement, and needs the source code of the program (1) to interpret it in the nonstandard way (2).
Notably, while the function (1) must be evaluated inside-to-outside (right-to-left), the chain rule (2) is an associative matrix product that can be evaluated in any order. This generally changes the number of operations and is a more open-ended computational problem. Forward-mode automatic differentiation seeds $\boldsymbol{J}_{1}=\boldsymbol{1}$ (the derivative of the input with respect to itself) and multiplies $\boldsymbol{J}_{N}(\boldsymbol{J}_{N-1}(\cdots(\boldsymbol{J}_{2}\boldsymbol{J}_{1})))$ by ?pushing? every column of $\boldsymbol{J}_{1}$ forward through the product in the same evaluation order as $\boldsymbol{f}$. Reverse-mode first computes $\boldsymbol{f}$ in a forward pass, then seeds $\boldsymbol{J}_{N}=\boldsymbol{1}$ (the derivative of the output with respect to itself) and multiplies $(((\boldsymbol{J}_{N}\boldsymbol{J}_{N-1})\cdots)\boldsymbol{J}_{2})\boldsymbol{J}_{1}$ by ?pulling? every row of $\boldsymbol{J}_{N}$ backwards through the product. This usually makes forward-mode faster when $\boldsymbol{f}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ has more outputs ($n\gg m$), and reverse-mode better when there are more inputs ($m\gg n$).
In practice, both modes are implemented using techniques called operator overloading or source code transformation. The former specializes every function on a particular ?dual number? type that propagates both the value and derivative of the function (e.g., Revels et al., 2016). The latter analyzes the source code for $\boldsymbol{f}$ and transforms it into another code that computes $\boldsymbol{J}$. In any case, automatic differentiation does not work on compiled binaries, but requires access to the source code to compute gradients with a different path through the instructions of the program.
@@
Figure 1: SymBoltz represents cosmological models with symbolic equations grouped in physical components for the metric, gravity, and particle species. This is compiled to a numerical problem that splits equations into background and perturbation stages and generates fast code for ODE functions $\boldsymbol{f}$ and Jacobians $\boldsymbol{J}$. The problem is then solved, and the result is stored in a solution object that gives access to any model variable $S$.
@@ 2 Code architecture and main features
SymBoltz is designed around three main features. In brief, it has a symbolic-numeric abstraction interface where users enter high-level symbolic equations that are automatically compiled to fast low-level numerical functions. It cures stiffness with implicit ODE solvers instead of approximation schemes to keep models simple, elegant, and extensible. It is differentiable, so it is possible to get accurate derivatives of any output with respect to any input. This provides rapid model prototyping, helpful abstractions, and automates tasks that must be done manually when modifying other codes. SymBoltz encourages interactive use and pursues a modular design that lets users integrate what they need from the package into their own applications. The end goal is to prioritize the modeler, who should be able to just write down their equations, while SymBoltz automates modeling chores.
The code is written in the Julia programming language (Bezanson et al., 2017), which has a rich ecosystem of scientific packages and aims to resolve the two-language problem. The symbolic-numeric interface is built on ModelingToolkit.jl [^9] https://github.com/SciML/ModelingToolkit.jl by Ma et al. (2022) and Symbolics.jl [^10] https://github.com/JuliaSymbolics/Symbolics.jl by Gowda et al. (2022). The compiled functions are solved by implicit ODE integrators in OrdinaryDiffEq.jl [^11] https://github.com/SciML/OrdinaryDiffEq.jl/ by Rackauckas and Nie (2017) and linear algebra methods in LinearSolve.jl [^12] https://github.com/SciML/LinearSolve.jl. Automatic differentiation works through ForwardDiff.jl [^13] https://github.com/JuliaDiff/ForwardDiff.jl by Revels et al. (2016).
The next subsections describe SymBoltz’ three main features in depth. Other implementation details are given in Appendix A. This paper describes SymBoltz version 1.0.0. We refer to the package documentation for definitive up-to-date information.
@@ 2.1 Symbolic-numeric interface
The core of SymBoltz is built around a symbolic-numeric interface illustrated in Fig. 1. A model defines what equations should be solved; a problem decides how they are solved; and a solution holds what was solved. Variables and equations are specified in a high-level user-friendly symbolic modeling language, then compiled to low-level numerical code that is integrated by ODE solvers. With knowledge of the symbolic equations, SymBoltz analyzes their structure programmatically to ease model specification, automates mechanical boilerplate tasks, and generates fast and stable code that avoids approximations. In our opinion, there are three properties that motivate such a symbolic abstraction layer around the Einstein–Boltzmann equations.
@@
First, the equations are most easily written as one large ODE, but optimally solved as several ODEs in stages such as the background, perturbations, and line-of-sight integration. Mathematically, the Jacobian matrix of the single ODE encodes which variables are independent and belong to which stages. It is easiest for modelers to specify equations in one common system, where the perturbations can refer to variables in the background, and then separate the computational stages automatically.
@@
Second, solving each stage is hard due to stiffness, performance requirements, and dependence on previous stages. To overcome this, it helps to automatically generate ODE code that is optimal in speed, accuracy, and stability, generate the ODE Jacobian needed by implicit solvers, and interpolate variables from previous stages. Modelers should not have to do this manually.
@@
Third, the equations are often modified, as we do not know the true cosmology. This can be made easier by organizing equations in a way that reflects the physical structure of the model, so it is easy to replace one related subset of equations with another without duplicating the entire model. This could be slow to do at the numerical level, but can be done with no runtime penalty in an intermediate symbolic representation of the equations.
@@
Inspired by this, SymBoltz inverts the traditional layout of Boltzmann codes. Most codes are built as pipelines following the background, perturbations, and other stages, but SymBoltz is primarily structured around the physical components that make up the equations. Related variables and equations are grouped in distinct components for the metric, gravity, photons, baryons, dark matter, dark energy, neutrinos, and other species (see Fig. 1). This isolates everything related to one submodel in one place. Any set of such submodels is joined into a full cosmological model, such as $\mathrm{\Lambda CDM}$. Interactions (e.g., Compton scattering or sourcing of gravity) are equations that connect components. Adding new submodels is easy, and SymBoltz is largely devoted to simply building a well-organized library of them.
@@
SymBoltz then takes a full cosmological model, separates it into stages, and generates numerical code to solve each stage. This preprocessing does not slow the code at runtime. To the contrary, the symbolic equations give extra information such as the Jacobian that enhances speed and stability. Modifications automatically enjoy these benefits without extra user input.
@@
The modular component-based structure scales well in model space. For example, modified gravity or dark energy models can be written as self-contained submodels and combined with other components to create many different extended cosmological models. The generic symbolic framework also handles reduced models with noninteracting radiation, matter and dark energy species. When exploring modified gravity theories, it can be helpful to test such toy models to understand how gravity responds to pure fluids without coupling to baryons and photons.
@@
In contrast, the monolithic layout of traditional codes scatters changes for one submodel across different modules for input, the background, perturbations, output, and so on. As more submodels are added, the code is increasingly intertwined and fragmented. They are also not flexible enough to test simpler toy models, as baryons and photons are inherently hardcoded. This design fits the computational procedure better than the physical structure. It can lead to an overwhelming code that is hard to read and modify. This complexity scales poorly in model space.
@@
As an alternative to component-based modeling, SymBoltz also includes a version of the default $\mathrm{\Lambda CDM}$ model with all equations packed into one big ?unstructured? system. This makes it trivial to change anything in the equations, but sacrifices modularity. The symbolic interface makes it very compact: SymBoltz defines a full $\mathrm{\Lambda CDM}$ model in only 277 lines of code, [^14] https://hersle.github.io/SymBoltz.jl/stable/LCDM/ while the equivalent code to browse in CLASS is spread over 10 files with 27721 lines.151515input.{h,c}, background.{h,c}, thermodynamics.{h,c}, perturbations.{h,c}, and wrap_recfast.{h,c} counted by wc.
@@
PyCosmo’s symbolic interface also generates code for ODEs and sparse Jacobians, but does not focus on component-based modeling or automating tasks such as stage separation. In this way, SymBoltz’ interface simplifies the solution and modification of the Einstein–Boltzmann equations. It aims to maximize speed, stability, and convenience with minimal user input. The next subsections explain how the symbolic features work in detail, and Sect. 3 shows them in action in a concrete example.
@@ 2.1.1 Automatic numerical code generation
SymBoltz automatically compiles symbolic equations to numerical code for ODEs
$\frac{\mathrm{d}\boldsymbol{u}}{\mathrm{d}\tau}=\boldsymbol{f}(\boldsymbol{u},\boldsymbol{p},\tau).$ (3)
The generated code is fast and prevents users unfamiliar with Julia or SymBoltz from writing slow code. If necessary, we can escape the standard code generation and call arbitrary numerical functions; for example, to solve a nonlinear equation for the minimum of a potential or interpolate tabulated data. The code generation handles tasks such as allocating indices for each ODE state $u_{i}$, and does optimizations such as common subexpression elimination (see Sect. 3.2). This is helpful as Boltzmann solvers tend to have large $\boldsymbol{f}$, and ODE solvers evaluate $\boldsymbol{f}$ many times.
@@ 2.1.2 Automatic handling of observed variables
A general ODE (3) has two types of variables: ?unknowns? $\boldsymbol{u}(\tau)$ are integrated with respect to time, while ?observed? variables are any functions of the unknowns. The Einstein–Boltzmann equations are usually written with many observed variables.
For example, consider the metric and gravity equations in Appendices A.1 and A.2 sourced by some known $\rho(\tau)$, $\delta\rho(\tau,k)$, and $\Pi(\tau,k)$. Here, $a(\tau)$ and $\Phi(\tau,k)$ are the only unknown (differential) variables that the ODE is integrated for, whereas $z(\tau)$, $\mathscr{H}(\tau)$, $H(\tau)$, and $\Psi(\tau,k)$ are all observed (algebraically) from the unknowns. Furthermore, it is not always straightforward to express observed derivatives such as $\mathscr{H}^{\prime}$ or $\Psi^{\prime}$, but SymBoltz automatically expands them using the definitions of $\mathscr{H}$ or $\Psi$. Of course, we can eliminate all observeds by explicitly inserting them into the equations for the unknowns. However, this reduces readability, as observed variables are helpful intermediate definitions that break up the equations, and we may want to extract them from the solution as well. Furthermore, modified models can change the sets of unknown and observed variables (e.g., modified gravity can change the constraint equation for $\Psi$ into a differential equation). It is easier to build models when variables are not hardcoded as either unknown or observed.
SymBoltz reads a full system of equations as input, such as that defined by the entirety of Appendix A, and automatically separates unknown and observed variables. After solving the ODE for its unknowns, SymBoltz can automatically recompute any observed variable from its expansion in unknowns.
The bottom line is that users can easily use any variable anywhere by just referring to it, whether it is unknown or observed. In other solvers it can be necessary to look up observed expressions and recompute them manually. This is tedious, error-prone, and can tempt users to unnecessarily make simplifying assumptions (e.g., approximating $\Psi\approx\Phi$ without anisotropic stress).
@@ 2.1.3 Automatic stage separation and splining of unknowns
In principle, the entire Einstein–Boltzmann system (i.e., background and perturbations) can be integrated all at once. However, it can be broken down into sequential computational stages that each depend only on those before it. To alleviate stiffness in each stage, avoid recomputing the background for every perturbation mode, and to improve performance by integrating smaller ODEs, all Boltzmann codes solve the system stage-by-stage and spline variables from one stage as input to the next.
@@
To illustrate this case, again consider the general relativistic equations in Appendix A.2 sourced by some known $\rho(\tau)$, $\delta\rho(\tau,k)$, and $\Pi(\tau,k)$. Clearly, $\Phi(\tau,k)$ and $\Psi(\tau,k)$ depend on $a(\tau)$, but $a(\tau)$ does not depend on $\Phi(\tau,k)$ or $\Psi(\tau,k)$, reflecting the perturbative nature of the problem. We can first solve for only $a(\tau)$ in the ?background,? then spline and look up $a(\tau)$ to solve for $\Phi(\tau,k)$ and $\Psi(\tau,k)$ in the ?perturbations,? instead of solving for all three together and repeatedly integrate $a(\tau)$ for every $k$.
@@
SymBoltz uses the same stage separation strategy as other Boltzmann codes, but automates it with knowledge of the symbolic equations. First, all equations are split into background and perturbation stages. A Cubic Hermite spline is then constructed for all background unknowns (e.g., $a(\tau)$, $X_{\mathrm{H}}^{+}(\tau)$, $X_{\mathrm{He}}^{+}(\tau)$, $T_{b}(\tau)$, and $\kappa(\tau)$) and replaces the background unknowns in the perturbations. One vector-valued spline is used to avoid repeatedly looking up $\tau$ for every unknown. Hermite splines are optimal for interpolating ODEs, as they take both $\boldsymbol{u}(\tau)$ and $\boldsymbol{u}^{\prime}(\tau)$ into account for better accuracy, and $\boldsymbol{u}^{\prime}(\tau)$ is known analytically from $\boldsymbol{u}(\tau)$ through the ODE (3). In contrast, observed variables (e.g., $\mathscr{H}(\tau)$) are computed from the (splined) unknowns, as their derivatives are not known directly and splining them is less accurate. This makes any background variable available in the perturbations ?for free.? All stages of the Einstein–Boltzmann equations are solved as if they are one common system of equations, while stage separation is done automatically under the hood.
@@
The separation into background and perturbation stages is guaranteed. It just reflects the perturbative structure of the linear Einstein–Boltzmann equations, where each order depends only on lower orders. However, they can often be broken further down: most thermodynamics (recombination) models can be separated from the background, and integral solutions (e.g., the optical depth $\kappa(\tau)=\int_{\tau_{0}}^{\tau}\kappa^{\prime}(\bar{\tau})\mathrm{d}\bar{\tau}$ and line-of-sight integration) can be computed after solving the differential equations. Future work could extend SymBoltz’ background-perturbations separation to split all finer stages by symbolically inspecting the dependencies between variables in the equations.